OpenStudy (anonymous):

find the area of the region that is bounded by these equations,, 1) y=lnx, x-axis, x= e^2 2) y = x^2, y = 8-x^, 4x - y +12 =0

OpenStudy (anonymous):

have you done integration yet?

OpenStudy (anonymous):

not yer

OpenStudy (anonymous):

is calculus class?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then I think only way to solve it is by using integration. Are you familiar with the concept of integration

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So graph two function in calculator. Which one is on top?

OpenStudy (anonymous):

you know how ln(x) curve look like. Also x=e^2 is verticle line.

OpenStudy (anonymous):

ln(x) is 0 when x=1 right? so you will \[\int\limits_{1}^{e^2}\ln(x) dx\]

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

the answer here in the handout is \[e ^2 +1\]

OpenStudy (anonymous):

i got it now... :)

OpenStudy (anonymous):

but how can i draw y=lnx and x= e^2 in the cartesian plane

OpenStudy (anonymous):

e^2 verticle line

OpenStudy (anonymous):

how can i solve for its x and y coordinate?